A body is dropped from a height H. The time taken to cover second half of the journey is

Grade:12

4 Answers

Apurva Sharma
5 years ago
the initial speed of the body,u=0
let ‘T’ be the total time taken by body to reach ground.
thus height covered:
H=1/2 gT2 
T=(2H/g)1/2
 
now let the time taken for 1st half of journey be ‘t’
distance covered th time t:
1/2H= 1/2gt2
t=(H/g)½
 
now time taken for second half of journey : T-t
= (2H/g)1/2 -(H/g)½
thus time taken for second half of journey is :  (H/g)1/2 * (2-1)1/2
= (H/g)1/2
ch vineeth
5 years ago
initial speed is zero hence u=0 and the distance to be travelled is H/2 . we know that in rectilenear motion of the body the relation between u s a and time t is s=ut +1/2(at2)     if we substitute all values we got in that equation we get h/2=1/2gt2 as the acceleration of the body is acceleration DUE TO GRAVITY hence we get t =(gh)½  …...this is the answer and the value of g is approximately 9.8        .
 
 
ch vineeth
5 years ago
answer is h/g½                                                                                                          
                                                                                                                                  
                                                 . 
Sheikh rafeya
5 years ago
We know that s=ut+1/2gt and initialvelocity is u=0and height isH/2 so on substituting the in the formula we get t = H/g. so this is the answer